OpenStudy (anonymous):

i need some help finding the antiderivative of 5x(x^2+1)^4

OpenStudy (dls):

So you just need to integrate it. \[\LARGE \int\limits 5x(x^2+1)^4 dx \] using by parts.

OpenStudy (nincompoop):

erm

OpenStudy (dls):

\[\Large \int\limits\limits 5x(x^2+1)^4 dx = 5x \int\limits (x^2+1)^4 - 5 \int\limits(x^2+1)^4dx\]

OpenStudy (dls):

Alternatively, \[\large (x^2+1)^4= 4C0 (x^2)4 + 4C1 (x^2)^3+4C2 (x^2)2 + 4C3 (x^2) + 4C4\]

OpenStudy (nincompoop):

wasteful use substitution

OpenStudy (dls):

oh yeh :P didn't strike :3 x^2+1=t woot

OpenStudy (nincompoop):

after factoring out 5 use u substitution and you are 2 steps away to completion

OpenStudy (dls):

he is ryt

OpenStudy (nincompoop):

\[5 \int\limits x(x^2+1)^4dx \] \[u = x^2+1;du = 2xdx\] \[\frac{5}{2} \int\limits u^4du\] do the rest

OpenStudy (anonymous):

i still dont get it

OpenStudy (larseighner):

This is a good time for advanced guessing. You see the exponent 4 on the quantity \( (x^2 + 1) \). So ask yourself what would the derivative of \( (x^2 + 1)^5 \) look like? That's a chain rule problem.

OpenStudy (larseighner):

The derivative of \((x^2 + 1)^5\) would be \( 5(x^2 +1)^4(2x) \) , right? That's just the chain rule. So in better order it would be \( 10x(x^2 +1)^4 \). Now compare that to your integrand.

OpenStudy (larseighner):

Whoa! The derivative of \((x^2 + 1)^5\) is exactly twice your integrand. So what would the derivative of \({1 \over 2}(x^2 + 1)^5\) be?

OpenStudy (nincompoop):

here, read up on substitution method https://www.math.hmc.edu/calculus/tutorials/substitution/

OpenStudy (nincompoop):

after completing the integration using the u-subtitution, you will need to return it back to what it was so you have an x and not U

OpenStudy (nincompoop):

there is no guess-work here, but just a use of the algebra you've learned your entire life and utilizing it more systematically

OpenStudy (larseighner):

Advanced guessing, of course, is not guess-work. It is a way to avoid needless substitution. The uniqueness theorem assures that if you come up with an antiderivative, it is the only answer, although it may not be in exactly the same form.

OpenStudy (anonymous):

if i had something like integral x*e^(x^2) ... can we solve it like this: u=e^(x^2) du=(2x*e^(x^2))

OpenStudy (anonymous):

what should i do next?

OpenStudy (anonymous):

just let u = x^2 , du =2xdx you have it becomes \[\dfrac{1}{2}\int e^udu\] It's easy now, right?

OpenStudy (anonymous):

where did the 1/2 come from? @OOOPS

OpenStudy (anonymous):

hey, du = 2x dx, --> xdx = 1/2 du,

OpenStudy (anonymous):

Your original problem is \[\int xe^{x^2}dx\] right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

u=e^(x^2) du=x dx

OpenStudy (anonymous):

If you let u = e^(x^2) , du = 2xe^(x^2) du the whole integral is \[\int \dfrac{1}{2}du =\dfrac{1}{2}u +C=\dfrac{1}{2}e^{x^2} +C\]

OpenStudy (anonymous):

got it?